OpenStudy (anonymous):

(z/4) + 7 >= -5 A. z <= 10 B. z >= -48 C. z >= 35/4 D. z <= -35/4 E. z > 65 F. z >= 0 Explain how you got your answer please.

OpenStudy (anonymous):

(z/4)>=-5-7 z>=-12*4 z>=-48

OpenStudy (anonymous):

Thank you so much!